OpenStudy (phebe):

I'M CONFUSED PLZZZ HELP K Solve: 3/(m-4)+1/(3(m-4))=6/m

OpenStudy (anonymous):

It's 9

OpenStudy (phebe):

u sure

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

wait

OpenStudy (phebe):

so ur not sure

OpenStudy (anonymous):

I have to check the work

OpenStudy (phebe):

ok thanx

OpenStudy (anonymous):

Because 10/ 3(m-4) = 6/ m

OpenStudy (phebe):

yes i think thas tru

OpenStudy (anonymous):

And m-9/ (m-4)m=0. So M = 9 for the solution.

OpenStudy (phebe):

so thas the answer

OpenStudy (anonymous):

Should be yes.

OpenStudy (phebe):

ok thanx very much nd fan me back k XDXDXD

OpenStudy (anonymous):

Aha fer sure.

OpenStudy (phebe):

thanx lol